# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v3

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast. To this end, several end-to-end learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach addressing this issue, additionally providing an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-CoSMo. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a small radiological task, PnP-CoSMo allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
@@ 1 Introduction
Through comprehensive experiments, we shed light on several properties of PnP-CoSMo such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility with a specific radiological task.
@@ 2.1 Reconstruction methods for accelerated MRI
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke (2016) introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. (2016) proposed adaptive weighting-based guided CS and Pingfan et al. (2018) used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods (Zhou and Zhou, 2020; Yang et al., 2020; Liu et al., 2021a, b; Feng et al., 2022; Lei et al., 2023), on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet (Lei et al., 2023) is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
@@ 3.1.3 Plug-and-play content consistency operator
where a content/style model decomposes into content and style, followed by a replacement of this corrupted content with high-quality content derived from and composing the improved image from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss in the following section the design and learning of the content/style model .
@@ 3.2.3 Network architecture and content capacity
where and are the height and width of the input image and () and () are the those of the corresponding content maps. This ratio is controlled via the number of up/downsampling blocks in the networks. Our notion of content capacity is closely related to the more general concept of the locality bias of image-to-image models (Richardson and Weiss, 2021). A content/style model with a high content capacity has a large content resolution, allowing it to learn a rich, highly expressive content representation that strongly influences the output’s structure. Consequently. the model’s style spaces are restricted to learn mostly low-level global features.
@@ 3.3 Iterative reconstruction based on content consistency
The core assumption of an idealized content/style model is that contents and are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error111If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal network design and training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. We refer to this gap between and as content discrepancy. Fig.3a shows content discrepancy as observed empirically. During reconstruction, the error in would limit the efficacy of the operator , affecting the reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section 3.2.2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content , but only to undersampled k-space measurements , we aim at solving the following minimization problem in CR
@@ 4.1 Experiment design
We conducted 4 sets of experiments to comprehensively evaluate PnP-CoSMo – (a) simulations, (b) benchmark against end-to-end methods on NYU DICOM data, (c) benchmark and ablation on clinical multi-coil raw data, and (d) radiological evaluation. Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case was considered as well.
4.1.1 Simulation
@@
4.1.2 Benchmark on the NYU DICOM dataset
In the NYU benchmark, we compared PnP-CoSMo against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-CoSMo, which requires only image-domain training data, can outperform methods that additionally require k-space training data. Furthermore, end-to-end methods are task-specific where the reference and target contrast are fixed – e.g. T1W and T2W, respectively. On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality. Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks, namely T1W-guided T2W reconstruction and T2W-guided T1W reconstruction. The NYU brain dataset (Knoll et al., 2020; Zbontar et al., 1811) includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts. Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al. (2022). A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at and added Gaussian noise of . The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50). Images from the former two splits were used to pre-train, fine-tune, and validate the content/style model. End-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits. The recon-val split was used to tune the PnP-CoSMo algorithm, and recon-test was the held-out test set. We benchmarked PnP-CoSMo against one well-known single-contrast network – MoDL (Aggarwal et al., 2018) – and two recent multi-contrast networks – MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023). These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance – e.g. for the T1W-guided T2W reconstruction task, in-distribution means that the end-to-end model was trained on this task, whereas out-of-distribution means the model was trained of the T2W-guided T1W reconstruction task.
@@ 4.1.2 Benchmark on the NYU DICOM dataset
4.1.3 Benchmark and ablation on the LUMC multi-coil dataset
@@
In our benchmark on clinical multi-coil data, the goal was to test PnP-CoSMo on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance. Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table 1 shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
@@
As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate the content/style models. The models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of =1.8-2. We further undersampled it retrospectively to higher accelerations by dropping subsets of the acquired lines. An overview of the data split is shown in Table 2. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
@@
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-CoSMo with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) (Ehrhardt and Betcke, 2016). As an unguided plug-and-play baseline, we used PnP-CNN (Ahmad et al., 2020), and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used (PROSIT) (Mattern et al., 2020), which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-CoSMo, we ablated PFT and CR to assess their contribution, and finally, as an upperbound for PnP-CoSMo representing zero content discrepancy, we disabled PFT and CR and used the ideal content . In both NYU and LUMC benchmarks, we used three perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality (Kastryulin et al., 2023). All three metrics are bounded in where represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
4.1.4 Radiological evaluation
@@
Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-CoSMo reconstructions. The study was conducted on a small sample of the LUMC-TRA recon-test set. With the help of a junior radiologist, we selected 3 cases with brain metastases. The evaluation was performed by a senior neuroradiologist and comprised two sets of criteria, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. A diagnostic-quality reconstruction was defined as one that scored at least ”fairly diagnostic” on the overall diagnostic quality criterion. We evaluated PnP-CoSMo and PnP-CNN at 4 clinically realistic accelerations of . In total, 9 images per patient, including the (42=8) reconstructions and the clinical ground truth (acquired at =2 and reconstructed by the vendor software), were presented to the senior neuroradiologist who scored each image individually, blinded to the reconstruction method and .

### v4

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast. To this end, several end-to-end learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach addressing this issue, additionally providing an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-CoSMo. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the public NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
@@ 1 Introduction
Through comprehensive experiments, we (a) shed light on several properties of PnP-CoSMo such as convergence and robustness, (b) demonstrate its equivalent performance and superior generalizability compared to state-of-the-art end-to-end models, (c) demonstrate its applicability on real-world raw data, and (d) evaluate its potential clinical utility with a preliminary radiological task.
@@ 2.1 Reconstruction methods for accelerated MRI
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke (2016) introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. (2016) proposed adaptive weighting-based guided CS and Song et al. (2020) used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods (Zhou and Zhou, 2020; Yang et al., 2020; Liu et al., 2021a, b; Feng et al., 2022; Lei et al., 2023), on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet (Lei et al., 2023) is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
@@ 3.1.3 Plug-and-play content consistency operator
where a content/style model decomposes into content and style, followed by a replacement of this corrupted content with high-quality content derived from and composing the improved image from it. The model is learned separately via an offline procedure and constitutes the core of our plug-and-play design. Before formally defining the operator and developing the reconstruction algorithm, we discuss in the following section the design and learning of the content/style model .
@@ 3.2.3 Network architecture and content capacity
where and are the height and width of the input image and () and () are those of the corresponding content maps. This ratio is controlled by changing the number of up/downsampling blocks in the networks. Our notion of content capacity is closely related to the more general concept of the locality bias of image-to-image models (Richardson and Weiss, 2021). A content/style model with a high content capacity has a large content resolution, allowing it to learn a rich, highly expressive content representation that strongly influences the output’s structure. Consequently, the model’s style spaces are restricted to learn mostly low-level global features. By finding the optimal content capacity for a given two-contrast image-domain dataset, one can quantify the amount of shared spatial information in this dataset, which will have consequences on the effectiveness of guidance in the reconstruction.
@@ 3.3 Iterative reconstruction based on content consistency
The core assumption of an idealized content/style model is that contents and are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error111If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal network design and training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. We refer to this gap between and as content discrepancy. Fig.3a shows content discrepancy as observed empirically. During reconstruction, the error in would limit the efficacy of the operator , affecting the reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section 3.2.2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content , but only to undersampled k-space measurements , we aim at solving the following minimization sub-problem in CR
@@ 4.1 Experiment design
We conducted four sets of experiments to comprehensively evaluate PnP-CoSMo – (a) analyses on simulated MRI images, (b) benchmark against end-to-end methods on the public NYU fastMRI DICOM data, (c) benchmark and ablation on clinical multi-coil raw data, and (d) a preliminary radiological evaluation. Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case of T2W-guided T1W reconstruction was considered as well.
4.1.1 Analyses on simulated MR data
@@
4.1.2 Benchmark on public NYU fastMRI DICOM data
In the NYU benchmark, we compared PnP-CoSMo against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-CoSMo, which requires only image-domain training data, can outperform end-to-end methods that additionally require k-space training data and are optimized to directly solve the reconstruction problem. Furthermore, end-to-end methods are task-specific models where the reference and target contrasts are fixed, e.g. T1W as reference and T2W as target. On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality. Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks, namely T1W-guided T2W reconstruction and T2W-guided T1W reconstruction.
The public NYU fastMRI brain dataset (Knoll et al., 2020; Zbontar et al., 2018) includes multi-coil raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts. Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al. (2022). A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at and added Gaussian noise of . The dataset was split into 3 folds at the subject-level for cross-validation, with two partitions representing training data and one representing the test data in each fold. For PnP-CoSMo, images from the training set were used to pre-train and fine-tune the content/style model, whereas the end-to-end methods were trained on paired examples composed of the reference image, the target contrast k-space, and its ground truth image. We benchmarked PnP-CoSMo against one well-known single-contrast network – MoDL (Aggarwal et al., 2018) – and two recent multi-contrast networks – MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023). These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance – e.g. for the T1W-guided T2W reconstruction task, in-distribution means that the end-to-end model was trained on this task, whereas out-of-distribution means the model was trained of the T2W-guided T1W reconstruction task.
@@ 4.1.2 Benchmark on public NYU fastMRI DICOM data
4.1.3 Benchmark and ablation on in-house LUMC multi-coil datasets
@@
In our benchmark on clinical multi-coil data, the goal was to test PnP-CoSMo on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, T1W/T2W images are often not fully subject-wise paired, e.g. due to missing acquisitions, representing a realistic case of data imbalance.
Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table 1 shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
@@
Each of the two LUMC datasets was split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate the content/style models. The models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of =1.8-2. We further undersampled it retrospectively to higher accelerations by dropping subsets of the acquired lines. An overview of the data split is shown in Table 2. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
@@
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-CoSMo with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) (Ehrhardt and Betcke, 2016). As an unguided plug-and-play baseline, we used PnP-CNN (Ahmad et al., 2020), and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used PROSIT (Mattern et al., 2020), which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-CoSMo, we ablated PFT and CR to assess their contribution, and finally, as an upperbound for PnP-CoSMo representing zero content discrepancy, we disabled PFT and CR and used the ideal content . In both NYU and LUMC benchmarks, we used three perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality (Kastryulin et al., 2023). All three metrics are bounded in where represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
4.1.4 Preliminary radiological evaluation
@@
Finally, as a supplement to the LUMC benchmark, we conducted a preliminary radiological evaluation with the goal of assessing the quality of PnP-CoSMo reconstructions in a concrete clinical context. The study was conducted on a small sample of the LUMC-TRA recon-test set. First, all 20 subjects from this set (see Table 2) were screened by a junior radiologist who selected three best cases with prominently visible brain metastases. The scoring was then performed independently by two senior neuroradiologists. The scoring system comprised two sets of criteria, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and tissue contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. Furthermore, we used PnP-CNN as a non-guided plug-and-play reconstruction baseline. We evaluated the two methods at four clinically realistic accelerations of , which are close to the routine clinical acceleration of =2 for this sequence. In total, nine images per patient, including the (24=8) reconstructions and the clinical ground truth (acquired at =2 and reconstructed by the vendor software), were presented to the two observers who scored each image individually, blinded to the reconstruction method and . The scores were then processed as follows. Given a reconstruction method and , the scores for each Likert-scale criterion were expressed as an aggregate by computing the mean across the three subjects and the two observers. The number of visible metastases was expressed as the fraction of total metastases across the three subjects. Moreover, the agreement between the two observers was measured using Spearman’s rank correlation coefficient, computed separately on the Likert-scale scores and the metastasis count.